### v2

# Transitioning from perturbative to non-perturbative splittings within QCD jets

Abstract
The study of Quantum Chromodynamics (QCD) at ultra-relativistic energies can be performed in a controlled environment through lepton-hadron deep inelastic scatterings. In such collisions, the high-energy partonic emissions that follow from the ejected hard partons are accurately described by perturbative QCD. However, the lower energy scales at which quarks and gluons experience colour confinement, i.e. hadronization mechanism, fall outside the validity regions for perturbative calculations, requiring phenomenological models tuned to data to describe it. As such, hadronization physics cannot be currently derived from first principles alone. Monte Carlo event generators are useful tools to describe these processes as they simulate both the perturbative and the non-perturbative interactions, with model-dependent energy scales that control parton dynamics. This work aims to use jets - experimental reconstructions of final-state particles likely to have a common partonic origin - to inspect this transition further. Although originally proposed to circumvent hadronization effects, we show that jets can be utilised as probes of non-perturbative phenomena via their substructure. The charge correlation ratio was recently shown to be sensitive to hadronization effects. Our work further improves this sensitivity to non-perturbative scales by introducing a new selection based on the relative placement of the resolved SoftDrop splitting within the clustering tree, defined as the unclustering that resolves the jet’s leading charged particles.
I I - Introduction
The concept of jets and their substructure dates back to the first evidence of quarks and gluons from collimated sprays of hadrons in annihilation experiments of electrons and positrons Riordan (1992); Ellis (2014); Brandelik et al. (1979). Since then, significant advances in fundamental QCD have been made due to copious production of jets in relativistic hadron-hadron colliders Ali and Kramer (2011). Many of the state-of-the-art Monte-Carlo (MC) models Bahr et al. (2008); Bothmann et al. (2019); Sjostrand et al. (2006); Bierlich et al. (2022) that generate full collision events rely on both a perturbative Quantum ChromoDynamics (pQCD) description of parton showers (including a resummation over all orders) and non-perturbative QCD (npQCD) implementation of hadronization and underlying event mechanisms. These parameterized implementations rely on fitting legacy $e+e-$ annihilation data to Deep Inelastic Scatterings (DIS) and, finally, recent data from hadron colliders Skands (2010); Aguilar et al. (2022).
Recent effort in the high-energy physics community has been in the area of developing novel experimental algorithms that translate a jet clustering tree to a theoretically motivated description of a parton shower Dasgupta and Salam (2001); Dasgupta et al. (2013); Larkoski et al. (2014); Mehtar-Tani et al. (2020). Measurements from the Relativistic Heavy-Ion Collider (RHIC) and from the Large Hadron Collider (LHC) have shown that jet substructure observables, when calculated via grooming techniques such as SoftDrop (SD) Sirunyan et al. (2018a, b); Aad et al. (2020a); Acharya et al. (2020); Aad et al. (2020b); Adam et al. (2020); Abdallah et al. (2021); Acharya et al. (2022); ALICE (2022a, b), allow one to fit pQCD calculations at the first hard splitting without the need for large hadronization corrections. Namely, the groomed momentum fraction ($z_{g}$), i.e. the fractional transverse momentum carried by the softest emitted parton, fits very well to the pQCD splitting functions when evaluated at the first SD splitting. For subsequent emissions along the higher energy (hard) branch of a jet shower Dreyer et al. (2018), the shape of the splitting function gradually changes from a steeply falling pQCD-like $1/z$ distribution to a flatter shape due to phase space restrictions. However, so as the impact of npQCD effects grows Robotkova (2022); Kunnawalkam Elayavalli (2022). Previous works Apolinário et al. (2021) demonstrated that SD groomed clustering trees significantly improve the correlation between the formation time of the parton shower emissions and the formation time of the first jet SD unclustering step, regardless of choice of jet clustering algorithm Apolinário et al. (2021). However, subsequent unclustering steps also observe their parton shower correlation worsen, although with a large dependence on the chosen algorithm.
The current state of SoftDrop substructure studies therefore requires the assessment of non-perturbative contributions to the jet evolution history, as they become more and more meaningful to the late-time jet dynamics. This calls for the exploration of experimentally robust jet substructure observables sensitive to hadronization physics. In Chien et al. (2022) the charge correlation ratio was proposed as a possible observable that is able to discriminate between two different classes of hadronization models: the Lund string model Andersson et al. (1983a), and the cluster fragmentation model Amati and Veneziano (1979). In this manuscript, we show that jet substructure selection can largely increase this sensitivity, allowing for more stringent tests to hadronization mechanisms. For that, this manuscript is organized as follows: in section II we will present an overview of the main hadronization models widely used by Monte Carlo hadronic event generators Pythia Bierlich et al. (2022) and Herwig-Bahr et al. (2008); the jet selection procedure to enhance hadronization effects, based on the resolved SoftDrop, will be presented in section III; results for the charge correlation ratio are discussed in section IV, with final conclusions in section V.
II II - Hadronization Models Overview
We generally expect that the evolution of a hard scattered parton ends with the parton attaining an on-shell mass for hadronization. A theoretical description of such process, i.e. the metamorphosis Ellis et al. (1976) of partons into hadrons, is unknown and is counted as one of the important unsolved problems in Physics. Hence, the treatment of hadronization has to reflect the non-perturbative conversion of the coloured, outgoing final-state partons into colourless hadrons through a combination of analytical results and different QCD-based phenomenological models. There are two main classes of hadronization models - the Lund string, implemented in Pythia, and the cluster fragmentation approaches, implemented in the Herwig and Sherpa Gleisberg et al. (2009) MC event generators.
The Lund String model Andersson et al. (1983a, b); Sjostrand (1984); Andersson (2023) is based on the assumption (supported by lattice QCD results Bali and Schilling (1992)) that the confining force field between a colour and an anticolour charge, as in a quark/antiquark colour-singlet $q\bar{q}$, can be accurately described as a flux tube with linearly-increasing potential energy with respect to the inter-charge distance $r$. This is accounted by the Cornell potential $V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+k\,r$, originally proposed to describe quarkonium masses, where the Coulomb term becomes negligible for large $r$. In the linear confinement framework, as the extremities of a $q\bar{q}$ dipole move apart from their common production vertex, the stretching of the colour flux tube can be modeled by a massless relativistic string with string tension $k\approx 1$ GeV/fm and no transverse degrees of freedom. Depending on the energies of the original pair, it may evolve into a stable so-called “yoyo” mode for smaller production energies, identified as mesons with flavour inferred from their quark content, or into a sequence of string fragmentations, for higher energies. For the case of a high-energy $q\bar{q}$, the potential energy of the string increases with parton separation until it is high enough for a new parton/antiparton pair to be formed. This leads the string to be broken up by the formation of new $q\bar{q}$ colour-singlet systems. After successive fragmentation, the energy scale reduces until on-shell hadrons ($p_{\text{hadron}}^{2}=m_{\text{hadron}}^{2}$) are produced. In the Lund String model, this is the point at which the fragmentation process is assumed to end. This picture is expanded to include gluons in Pythia, allowing $q\bar{q}$ strings to stretch through a finite number of gluons.
The cluster hadronization framework is based on the colour “preconfinement” property Amati and Veneziano (1979). Partons perturbatively evolve from the hard scattering scale $Q$ down to $Q_{0}$, where quarks and gluons rearrange themselves into on-shell lumps of colour singlets - clusters - whilst not yet allowing on-shell hadrons to emerge. This intermediate stage of jet development is what is called pre-confinement. It is then assumed that confinement, describing the processes that bring the scale $Q_{0}$ down to hadron masses $O\left(m_{\pi},m_{p}\right)$, will convert these limited mass, colourless clusters into hadrons. This is where the cluster fragmentation model comes into play. Herwig’s implementation starts by splitting the gluons at the end of the parton shower into (di)quark/anti(di)quark pairs Bahr et al. (2008), constrained by the reminiscent decay phase-space. After this first step, the event is fully reduced to colour-connected quark/antiquark or diquark/antidiquark pairs, to be grouped in the model’s colour singlet clusters. The cluster mass spectrum is independent on the collision system and centre-of-mass energy and is peaked at low masses, meaning clusters can be taken as highly-excited hadron resonances and decayed, according to phase-space, into the observed hadrons. Nonetheless, the hadrons in the large-mass tail of the cluster spectrum require a cluster fission process to split them into lighter clusters before decaying. This is done by producing a $q\bar{q}$ pair from vacuum, forming two new clusters, each with one of the original partons. Once all clusters pass the model’s dynamical mass cutoff, the new set of clusters are finally decayed into pairs of on-shell hadrons.
III III - Event generation and jet substructure
DIS events provide a clean environment to study the hadronization mechanism and confinement dynamics, as they produce final states with typically low multiplicities, allowing for more precise measurements. DIS data is also required for the testing and calibration of the theoretical tools, such as MC event generators, used to study the quark-gluon plasma produced in heavy-ion collisions. Here, we build on recent work Chien et al. (2022) related to electric charge and formation time studies in DIS systems towards refining the sensitivity of novel experimental observables to hadronization effects. As such, we make use of both Pythia 8.306 and Herwig 7 to test a variety of jet observables and selections across the two different phenomenological models - Lund and cluster, respectively - described in the previous Section. We aim to inform future experimental measurements at the upcoming Electron Ion Collider (EIC), planned to start its activity by the start of the new decade.
As such, we simulate 40 million $ep$ DIS collisions using both MC event generators, anticipated to be 4% of the total integrated luminosity at EIC Chien et al. (2022). We impose the maximum EIC beam energies of $\sqrt{s_{e}}=18$ GeV and $\sqrt{s_{p}}=275$ GeV for the electron and proton beams, respectively. A minimum momentum transfer of $Q^{2}>50$ GeV2/c2 is applied on the hard scatterings to ensure the production of high $p_{T}$ jets and particles are selected with transverse momentum $p_{T}>200$ MeV/c to remove high-collinearity to the beam.
From the list of final-state particles, jet finding via the anti-kT algorithm Cacciari et al. (2008) as implemented in fastjet 3.3.0 Cacciari et al. (2012) with a resolution parameter of $R=1$ is performed and inclusive $p_{T,jet}>7$ GeV/c [^1] While we closely followed the procedure depicted in Chien et al. (2022), we focused on $p_{T,jet}>7~{}\rm{GeV/c}$ instead of $p_{T,jet}>5~{}\rm{GeV/c}$ to minimize the differences between the hadronization models. transverse momentum jets are selected within a pseudo-rapidity range of $-3.5<\eta_{jet}<1.5$. We then perform the SD grooming on jets via the Cambridge/Aachen (C/A) algorithm Dokshitzer et al. (1997) for geometric substructure analysis, since it enforces a strict angular ordering in the clustering trees. Accordingly, the first declustering step of these trees corresponds to the largest opening angle. For formation time studies, the SD algorithm is applied to $\tau$ re-clustered trees Apolinário et al. (2021, 2024) since the $\tau$ algorithm’s distance measure becomes a proxy for the (inverse) formation time of the clustered particles in the high-energy, soft and collinear limits. As such, the first declustering will generally have the shortest formation time value.
Figure 1: Top panel: particle transverse momentum spectrum for particles with $p_{T}>200$ MeV/c for Pythia’s simple shower in full red circles and Herwig’s dipole shower in open red circles; Bottom panel: jet transverse momentum spectrum with a $p_{T,jet}>7$ GeV/c selection and $-3.5<\eta_{jet}<1.5$ (same colour scheme); both panels have below them the ratios of Pythia spectra with respect to Herwig.
With the goal of comparing hadronization models in mind, we select the parton shower descriptions in Pythia and Herwig that produce the best matches between profiles of the event’s particles and jets, with respect to observables like azimuthal angle ($\phi$), pseudorapidity ($\eta$) and transverse momentum ($p_{T}$). For the chosen collision settings, the results matched the most for Pythia’s simple shower and Herwig’s dipole shower. For reference, Fig. 1 shows the particle (top) and jet (bottom) transverse momentum spectra for the DIS events, in full circles for Pythia simulations and in open circles for Herwig with the settings and selections mentioned above, along with the ratios with respect to Herwig results. We can see in the jet spectra that variations between the Monte Carlo generators fall below 10% for a wide range of transverse momenta.
The momentum fraction at each splitting along the primary branch of the clustering tree is defined as
$z_{g}=\frac{p_{T,2}}{p_{T,1}+p_{T,2}}\quad,$ (1)
where the transverse momentum of two objects $(1,2)$ are numbered such that $p_{T,2}\leq p_{T,1}$. The SD criterion imposed on the C/A re-clustered trees follows the general selection
$z_{g}>z_{cut}\left(\frac{\Delta R_{12}}{R}\right)^{\beta}\quad,$ (2)
where we set the soft threshold $z_{cut}=0.1$ to remove the divergence at small $z_{g}$ and the angular exponent $\beta=0$. The opening angle, commonly referred to as the groomed jet radius or $R_{g}$, is calculated as the distance between the two prongs in the rapidity($y$)-azimuthal($\phi$) angle space $\Delta R$, given by
$\Delta R_{12}^{2}=\sqrt{\left(y_{1}-y_{2}\right)^{2}+\left(\phi_{1}-\phi_{2}\right)^{2}}\quad.$ (3)
Finally, we can also evaluate the formation time $\tau_{\text{form}}$ Apolinário et al. (2024), which is the time that a quantum state, such as a parton, takes to behave as two independent sources of additional radiation, for example, the two daughter-partons from a splitting. The formation time within jets is associated to varying stages of the parton evolution, ranging from pQCD to npQCD effects. For a jet splitting, $\tau_{\text{form}}$ can be calculated by:
$\tau_{\text{form}}=\frac{1}{2\;E\;z\;(1-z)\;\left(1-\cos{\theta_{1,2}}\right)}\quad,$ (4)
where $E=E_{1}+E_{2}$ is the total energy of the two objects, $\theta_{12}$ the euclidean opening angle between the daughters’ 3-momenta and $z$ is the energy fraction of the softer daughter.
The definitions of $z_{g}$, $R_{g}$ and $\tau_{form}$ usually apply to the first SD jet clustering step from reclustered trees, but can be also used for any two quantities such as charged particles or other clusterings within the jet, as we will explore in this study.
Figure 2: Graphical representation of the clustering tree from a jet and the respective leading charged particles (LCP) in red circles, resolved SoftDrop splitting (RSD) in the red splitting and first SoftDrop splitting (1SD) in the blue splitting.
In this paper, we define three distinct splittings associated with the jet, diagrammatically represented in Fig. 2. The first unclustering step, starting from the fully clustered C/A jet, that passes our grooming criterion - first SoftDrop splitting (1SD, Fig. 2, blue unclustering) - will be the one we expect to follow a pQCD prescription, as it is the first splitting already shown to yield the Dokshitzer–Gribov–Lipatov–Altarelli–Parisi (DGLAP) splitting functions Larkoski et al. (2015) from perturbation theory.
The next we consider are the leading and sub-leading (in terms of their $p_{T}$) charged particles (LCP, Fig. 2, red circles) within the jet. Denoting these two objects as $1,2$ in Eqs. (1), (3) and (4) we can identify the parent particle. This is not a realistic splitting in a theoretical sense, since charged particles are produced post hadronization. However, it serves as an useful jet analysis tool, as the LCP is independent of the re-clustering procedure and can be directly measured in experiment. This was previously introduced in Chien et al. (2022).
Lastly, as a proxy for this splitting within the jet shower, we define the resolved SD splitting (RSD, Fig. 2, red unclustering) which is the first SD unclustering along the primary branch of the jet tree wherein the leading and sub-leading charged particles are separated into their own independent prongs. This clustering step refers to a specific point along the tree where the high-energy, hadronized remnants of the jet are resolved, providing a handle on the depth into the fragmentation pattern at which the two hardest final-state objects from the jet were produced. This can be done via split counting, giving us the so-called RSD depth $N_{RSD}/N_{SD}$, with $N_{RSD}$ accounting for the number of SD splittings along the main jet branch up to and including the RSD, while $N_{SD}$ accounts for the total number of SD splittings in the main branch.
Since vacuum parton showers typically generate angular ordered emissions and the re-clustering procedure is performed with an angular ordering variable via the Cambridge/Aachen algorithm, the first unclustering step to survive the SD criterion will produce prongs with larger $R_{g}$ within the jet (proxy for the opening angle), followed by a sequential reduction in the angle as we travel along the clustering tree. This is reflected on the top panel from Fig. 3, where the 1SD, in the blue markers (full circles for Pythia simulations and open circles for Herwig), will be on average observed at larger angles ($R_{g,1SD}\approx 0.52/0.56$ for Pythia/Herwig). Since the RSD, by construction, happens either at or after the 1SD in the clustering tree, their opening angles in the black markers are shown to be smaller ($R_{g,RSD}\approx 0.33/0.40$). As for the LCP, represented via the blue markers, we find that it closely follows the $R_{g}$ distribution of the RSD splitting, meaning the subsequent branching of the RSD daughters will mostly preserve the opening angle between the leading charged particles ($R_{g,LCP}\approx 0.33/0.40$). The $R_{g}$ is shown to be sensitive to the hadronization procedure, with significant differences between Pythia and Herwig.
Figure 3: Top panel: groomed opening angle distributions; Bottom panel: groomed momentum fraction distributions; the 1SD, RSD and LCP distributions are self-normalized and represented in red, black and blue markers, respectively, in full circles for Pythia jets and open circles for Herwig jets.
The pQCD asymmetry in the transverse momentum sharing expected for early splittings (splitting functions highly-peaked at small $z_{g}$) can also be observed in Fig. 3, in the bottom panel, with the 1SD being the most asymmetric splitting, peaking at $z_{g}=z_{cut}=0.1$, followed by the RSD and lastly the LCP, both tendentiously symmetrical. The results for both Monte Carlo generators are in good agreement with each other.
IV IV - Formation Time and Charge Correlation Ratio
The formation time distributions, calculated according to Eq. (4) for the 1SD, RSD and LCP splittings from the jets, are shown in Fig. 4, along with the ratios with respect to a LCP baseline on the bottom. The 1SD will naturally yield smaller $\tau_{form}$ values as they are selected at the earliest unclustering step, with the largest opening angles and smallest groomed momentum fractions. The LCP $\tau_{form}$ has a characteristic tail at large formation time since the two leading charged particles will predominantly have smaller angles as compared to the jet splittings. We recall that the RSD marks the splitting where the two leading charged particles follow different clustering steps. Interestingly, the RSD distribution is very close to the 1SD for early times ($\tau_{form}\lesssim 1~{}\rm{fm/c}$, around the proton radius), while its shape is identical to the LCP for later times, as the black plateau in the ratio panel from Fig. 4 shows. We note that the transition seems to occur around $\tau_{form}\in[1,5]~{}\rm{fm/c}$.
Figure 4: Formation time distributions for the 1SD (red), RSD (black) and LCP (blue) splittings from the sampled jets, where full circles are used for Pythia ratios and open circles for Herwig.
With RSD focusing on the high-energy charged particles within the jet, we can directly compare this study with the charge-correlation ratio observable introduced here Chien et al. (2022). This is defined as the ratio of yields of equally-charged, $d\sigma_{hh}/dX$, and oppositely-charged, $d\sigma_{h\overline{h}}/dX$, leading charged hadrons of flavour $h=\pi^{+},K,p$ (pions, kaons and protons), within the jets
$r_{c}(X)=\frac{d\sigma_{hh}/dX-d\sigma_{h\overline{h}}/dX}{d\sigma_{hh}/dX+d\sigma_{h\overline{h}}/dX}\quad,$ (5)
with respect to a given kinematic variable $X$.
Therefore, $r_{c}$ acts as a measure of the probability of producing jets with a given LCP charge configuration, with values ranging between $11\leq r_{c}\leq-$ 1: positive $r_{c}$ means a greater probability of jets having both positive or both negative LCP (same sign), while negative $r_{c}$ signals that jets tend to have both a positive and a negative leading charges (opposite sign). This was recently shown as a direct test of the string breaking phenomenon.
Since the definition of the $r_{c}$ is meant to be sensitive to hadronization via the two hardest hadrons, we can study it as a function of the formation time defined in Eq. (4). The charge ratio dependence of the LCP formation time for jets with leading pions, kaons and protons is shown in Fig. 5  Chien et al. (2022), in the black, red and blue markers, respectively, sampled from jets with at least two identified charged particles. The analysis of Pythia(Herwig) jets, corresponding to the charge ratios in full(open) circle markers, reveals that 55%(57%) have a leading pair of charged pions, 5%(3%) have leading kaons and 4%(6%) have leading protons.
Figure 5: Charge ratios with respect to the LCP formation time for jets with leading charged pions (black), kaons (red) and protons (blue), where full circles are used for Pythia ratios and open circles for Herwig.
Firstly, the charge ratios are shown to be overall negative, consistently across hadron species. However, the $r_{c}$ is clearly dependent on the formation time of the LCP. For large $\tau_{\text{form, LCP}}$, the charge ratios are approximately constant at large negative values of $r_{c}$, particularly for $KK$ and $pp$ pairs within jets. In fact, jets with leading protons have a near 100% probability of having $p\bar{p}$, with highly unlikely $pp$ and $\bar{p}\bar{p}$ configurations. This reveals a strong preference towards the production of opposite leading charged particles for late LCP jets. We recall that RSD is a good proxy for the LCP for large $\tau_{\text{form, LCP}}$ and that $\tau$ re-clustered trees will have the largest formation time values for the latest unclustering steps, i.e., in the last stages of the jet fragmentation pattern. As such, large LCP formation time selects jets whose leading charged hadrons get resolved very late, making them less likely to have subsequent branching randomizing the correlation between the charges.
However, for small $\tau_{\text{form, LCP}}$, the charge ratios become significantly closer to zero, indicating a more equitable probability of having same and opposite sign LCP jets. This result suggests a randomization of the LCP charge profile for early LCP jets, for whom the leading charged particles are resolved early on the jet fragmentation pattern, with several following unclustering steps that smear the charge correlation. In this formation time region, roughly around $\tau_{\text{form, LCP}}\lesssim 5$, Pythia noticeably diverges from the Herwig predictions for the charge ratio. Therefore, we identify a transition region $\left(1\lesssim\tau_{\text{form, LCP}}\lesssim 5\right)$ fm/c, that matches the $\tau_{form}$ distribution (Fig.4). The transition region is shaded in grey in both Figs. 4 and 5.
IV.1 Resolved SoftDrop depth
Our goal is to increase the sensitivity of this charge ratio variable to hadronization physics. Namely, jets can have their resolved SD earlier or later in their respective fragmentation pattern. When selecting jets with an identifiable RSD splitting that resolves the leading charged particles from that jet, we observe $\langle N_{SD}\rangle\approx 3$ SoftDrop unclusterings along the main branch for both for Pythia and Herwig, as can be seen in Fig. 6 on the top and bottom panels, respectively. As such, the contributions coming from jets with RSD in the first, second and third SDs are shown as stacked $N_{SD}$ distributions in red, blue and green bars, respectively. We find that 52% of Pythia jets and 59% of Herwig jets have their resolved splitting at the very first SD unclustering of the clustering tree. The RSD will fall on the second SD for an additional 30%/26% of jets and on the third SD for 13%/11% of them for the kinematic range studied.
Figure 6: Self-normalized distributions of the total number of SD splittings along the main jet branch (black markers), with the contributions from jets with their first (red), second (blue) and third (green) splittings resolving the leading charged particles (top panel for Pythia and bottom panel for Herwig).
Since jets will have a varying number of total SD splittings, we additionally show the RSD depth $N_{RSD}/N_{SD}$ in Fig. 7 with the respective contributions from the location of the RSD along the primary branch (same colour scheme). Jets with an RSD = 1SD that are located at $N_{RSD}/N_{SD}\approx 1$ are jets with only one SD emission along the branch and, although present, are typically rare. Jets with an RSD = 1SD will generally have an early RSD depth, and its distribution will be mostly related to the total number of SD emissions along the primary branch. It is clear that around $N_{RSD}/N_{SD}\approx 0.5$ the population of jets with an early RSD (RSD = 1SD) is no longer the dominant one. As such, we would expect that for this jet population the hadronization effects on the leading charged particles would become more visible. Considering that the charge dipole ratio presents a sensitivity to hadronization effects as well, we proceed to analyze how the $r_{c}$ can possibly depend on the RSD depth.
Figure 7: Top panels: self-normalized distributions of the depth of the resolved splitting within the clustering tree (black markers), with the contributions from jets with their first (red), second (blue) and third (green) splittings resolving the leading charged particles (top panel for Pythia and bottom panel for Herwig).
As such, for completeness, in Fig. 8, we further show how the $r_{c}$ depends on jet substructure topology. Namely, we plot the charge ratio as a function of the RSD depth for both Pythia (full circles) and Herwig (open circles). For both models, the earlier the RSD takes place, the more random will be the hadronization of the leading charged particles. As the RSD depth increases, the $r_{c}$ becomes more negative, highlighting that jets with late-stage RSD indeed strongly preserve an opposite charge relation for their LCP.
Figure 8: Charge ratios with respect to the RSD depth for jets with any leading hadronic flavours, where full circles are used for Pythia ratios and open circles for Herwig.
IV.2 Jet selection based on RSD
Having clearly identified a dependence of the charge correlation ratio on jet substructure, we will use the additional information provided by the RSD depth to further increase the differences between the hadronization models. Namely, from Fig. 7, it is possible to identify that jets in which the leading charged particles are coming from clustering steps further along the jet clustering tree will have typically its RSD depth $N_{RSD}/N_{SD}>0.5$. Using this observation, Fig. 9 replicates the results from Fig. 5, but this time selecting separately the jets with early RSD - $N_{RSD}/N_{SD}\leq 0.5$ shown on the top panel, and jets with late RSD - $N_{RSD}/N_{SD}>0.5$ shown on the bottom panel. The $N_{RSD}/N_{SD}\leq 0.5$ jet selection maintains a similar qualitative behaviour between Pythia’s and Herwig’s $r_{c}$ dependence on formation time, with an overall strengthening of the LCP opposite sign correlation for later time. However, the $N_{RSD}/N_{SD}>0.5$ selection showcases significant discrepancies between the generators for small $\tau_{\text{form, LCP}}$ jets. The charge ratios computed with Pythia do not have a significant dependence on $\tau_{\text{form, LCP}}$, regardless of the hadronic species of the LCP, leading to approximately constant pion, kaon and proton $r_{c}$, as evident on the bottom panel of Fig. 9. However, for the Herwig charge ratios, while late time still verifies a plateau-like behaviour, there is a steep climb towards $r_{c}=0$ for early-time LCP jets, significantly diverging from Pythia’s behaviour.
Figure 9: Charge ratios with respect to the LCP formation time for jets with leading pions (black), kaons (red) and protons (blue), where full circles are used for Pythia ratios and open circles for Herwig; top panel shows early-RSD jets, with small depths; bottom panel shows late-RSD jets, with large depths.
Figure 10: Charge ratio with respect to the LCP formation time for jets with late-RSD and with leading charged pions (black), kaons (red) and protons (blue); full markers belong to jets obtained with the DIRE parton shower model and open markers for VINCIA, both from Pythia 8.306; bottom panel shows the ratio with respect to the $r_{c}$ from Pythia’s simple model.
Since we are interested in the charge ratio’s sensitivity to hadronization features and selecting jets based on jet substructure variables, we need to check the resilience of our findings against pQCD effects, which can be done by comparing the charge ratios obtained with different parton shower descriptions provided by the two generators. Fig. 10 shows Pythia’s charge ratios for $N_{RSD}/N_{SD}>0.5$ jets and Fig. 11 the $r_{c}$ from Herwig, with all other parton shower models available in these Monte Carlo. Therefore, Fig. 10 has the $r_{c}$ from Pythia simulations using the DIRE (full markers) and VINCIA (open markers) showers and Fig. 10 the $r_{c}$ from Herwig simulations with the default shower. The panels below have the ratios calculated with respect to the simple and dipole baselines we have been using for Pythia and Herwig, respectively. The plateau nature of the Pythia $r_{c}$ is shown to be independent of the parton shower model employed, with variations limited to less than 10% relative to the simple shower baseline. In particular, the variations of the $r_{c}$ for jets with leading kaons (red) and leading protons (blue) are smaller than 2%. For the Herwig charge ratios, we observe that there is only a slight dependence on parton shower models of the early-time $r_{c}$, where jets with leading charged pions have a slightly less random charge configuration, jets with leading kaons have a stronger correlation and jets with leading protons remain basically the same. We can therefore conclude these charge ratio results to be robust against changes in the perturbative description, isolating the $r_{c}$ sensitivity to npQCD effects.
Figure 11: Charge ratio with respect to the LCP formation time for jets with late-RSD and with leading charged pions (black), kaons (red) and protons (blue); results obtained with the default parton shower model from Herwig 7; bottom panel shows the ratio with respect to the $r_{c}$ from Herwig’s dipole model.
V V - Conclusions
Jet substructure tools aimed to reduce non-perturbative contributions unlock a direct comparison between experimental results and analytical calculations. In this manuscript, by excluding the jet population in which non-perturbative contributions are reduced, we identified a jet sample where hadronization models yield qualitatively different results for the charged correlation ratio. By introducing the resolved SoftDrop splitting, as the location in the jet clustering tree in which the leading and sub-leading charged particles within a jet are resolved in independent branches, we show that the depth at which this splitting is located dictates the overall charge correlation ratio behaviour of the two hadronization models used in this work: Lund strings via Pythia and cluster fragmentation via Herwig.
We know from Fig. 1 that the overall event distributions from Pythia and Herwig are not exactly the same, conflating this comparative exercise with other factors. However, the shear magnitude of the differences observed for the two generators in the late-RSD ($N_{RSD}/N_{SD}>0.5$), early-LCP ($\tau_{\text{form, LCP}}\lesssim 1$ fm/c) jet selections lead us to conclude the differences observed come predominantly from the different hadronization models - Lund string and cluster fragmentation. This conclusion is reinforced by the robustness of the $r_{c}$ against varying parton shower descriptions. As such, this selection allows to highlight discrepancies coming from the string model, which retains memory of the incoming parton on the LCP fragmentation, as opposed to the cluster model that lumps the partons in clusters to achieve colour neutrality. In this case, for small $\tau_{form,LCP}$, one expects subsequent unclustering steps to randomize the resulting hadronization pattern.
Therefore, a selection based on the RSD depth reveals a qualitatively different behaviour between the charge ratios predicted by Pythia and Herwig, magnifying the sensitivity of this variable to the hadronization physics described by the two models.
VI Acknowledgments
We thank Isaac Mooney, Laura Havener, Yang-Ting Chien, Youqi Song, Krishna Rajagopal and Jose Guilherme Milhano for their valuable discussions and comments on the manuscript. This work was supported by the European Council (ERC) project ERC-2018-ADG-835105 YoctoLHC; by OE Portugal, Fundação para a Ciência e a Tecnologia (FCT), I.P., project CERN/FISPAR/0032/2021. LA and NOM acknowledges the financial support by FCT under contract 2021.03209.CEECIND and under PhD grant PRT/BD/154611/2022, respectively. RKE would also like to acknowledge funding by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grant number DE-SC0024660.